# How to Adapt Control Barrier Functions? A Learning-Based Approach with Applications to a VTOL Quadplane

### v3

@@ I INTRODUCTION
Figure 1: Conceptual illustration of the inner safe set and locally validated CBF parameter. (a) Candidate inner safe set $\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ defined via an ICCBF with a CBF parameter $\tilde{\bm{\alpha}}$. At boundary point ${\bm{x}}_{1}$, the CBF-constrained dynamics $G({\bm{x}}_{1};\tilde{\bm{\alpha}})$ is non-empty, whereas at ${\bm{x}}_{2}$ it is empty, indicating that the set cannot be rendered forward invariant. (b) With locally validated CBF parameters, the trajectory remains within the inner safe set shown in (a) over the finite-horizon, ensuring safety for that interval. By adapting the CBF parameters, the corresponding inner safe set is reshaped dynamically, alleviating conservatism by allowing the trajectory to extend beyond a fixed, globally verified inner safe set $\mathcal{C}^{*}(\bm{\alpha})$.
@@ II-A Tangent Cone
[…]
@@ III METHOD OVERVIEW
There exist values of $\tilde{\bm{\alpha}}$ for which, at some states ${\bm{x}}\in\partial\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ (see Fig. 1a), the set $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}})$ is empty, resulting the corresponding set $G({\bm{x}};\tilde{\bm{\alpha}})$ is also empty. This implies that the inner safe set $\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ cannot be rendered forward invariant by any control law satisfying the candidate ICCBF constraint.
@@
Consider the closed-loop system (2) with control inputs ${\bm{u}}\in K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}}(t))$. Suppose that the system is initialized with a locally validated CBF parameter $\tilde{\bm{\alpha}}(t_{0})$ at $t=t_{0}$, and that it is adapted at times $(t_{k})_{k\in\mathbb{N}\cup\{0\}}$ with $t_{k+1}-t_{k}<T$ for all $k$ such that $\tilde{\bm{\alpha}}(t_{k})$ satisfies (22) over $[t_{k},t_{k}+T]$. Then, the system trajectories remain within $\mathcal{S}$ for all $t\geq t_{0}$.
@@
We prove this by showing that for all $k\in\mathbb{N}\cup\{0\}$, the state remains in $\mathcal{S}$ over the interval $[t_{k},t_{k}+T]$. The proof is by induction.
@@
Inductive Step: Suppose the claim is true for some $k>0$, i.e., the state remains in $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{k}))\subseteq\mathcal{S}$ for all $t\in[t_{k},t_{k}+T]$. By the adaptation, the CBF parameter is updated at some time $t_{k+1}$ such that it is locally validated over $[t_{k+1},t_{k+1}+T]$, where $t_{k+1}<t_{k}+T$. Again, by Definition 7 and Nagumo’s Theorem, the system is ensured to remain in $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{k+1}))\subseteq\mathcal{S}$ over $[t_{k+1},t_{k+1}+T]$. ∎
@@ V-A Probabilistic Ensemble Neural Network Model
We employ a Probabilistic Ensemble Neural Network (PENN) as the prediction model, denoted as ${\bf F}$, to capture both prediction uncertainty and inherent output variability [18]. Building on the input $\bar{X}$ from Sec. IV, we augment it with the CBF parameter of interest $\tilde{\bm{\alpha}}$. Given this augmented input $X=[\bar{X}^{\top},\tilde{\bm{\alpha}}]^{\top}$, the model outputs two quantities $Y=[\phi,\delta]^{\top}$: a safety margin $\phi\in\mathbb{R}$ and a performance metric $\delta\in\mathbb{R}$ (see Fig. 2b). Concretely, we define $\phi$ as the minimum of the ICCBF constraint function over the interval $[t,t+T]$:
$\phi\coloneqq\min_{\tau\in[0,T]}\,b_{r}\left({\bm{x}}_{t+\tau|t},\,{\bm{u}}_{t+\tau|t};\,\tilde{\bm{\alpha}}\right).$ (24)
Notably, this minimum value is sufficient to verify the local validity condition because, as stated in Definition 7, the condition (22) is only non-trivial on the boundary of the inner safe set $\partial\mathcal{C}^{*}(\tilde{\bm{\alpha}})$. If $\phi$ remains above zero, then $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}})$ is non-empty throughout $[t,t+T]$, ensuring that the system stays within $\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ over that interval. Meanwhile, $\delta$ reflects the controller’s performance under the CBF parameter $\tilde{\bm{\alpha}}$ (e.g., capturing the rate at which the robot progresses toward a goal); we will discuss its criteria and usage in Sec. V-C.
@@ V-B2 Verification Under Aleatoric Uncertainty
$\textnormal{CVaR}_{\varepsilon}^{{\bf E}_{b}(X)}\left(\hat{\phi}\right)\coloneqq\sup_{\eta\in\mathbb{R}}\left\{\eta-\frac{1}{\varepsilon}\mathbb{E}_{{\bf E}_{b}(X)}\left[(\eta-\hat{\phi})^{+}\right]\right\},$ (27)
@@ V-C Online CBF Parameter Adaptation
From the verified set $\mathcal{A}$ (30), we select the final CBF parameter based on a performance metric $\delta\in Y$. Here, $\delta$ is defined as the accumulated time over the forward propagation interval $[t,t+T]$ during which the system’s speed towards the goal falls below a specified threshold, with lower values indicating more efficient progress. Hence, we adapt the CBF parameters by choosing: $\tilde{\bm{\alpha}}^{*}=\operatorname*{arg\,min}_{\tilde{\bm{\alpha}}\in\mathcal{A}}\hat{\delta}$, where $\hat{\delta}$ is the predicted performance metric associated with the candidate parameter $\tilde{\bm{\alpha}}$. This yields a less restrictive CBF-constrained control input set $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}}^{*})$, thereby improving performance.
@@ VI-A1 Aerodynamics Forces
Let $(u,w)$ be the velocity in the body frame, i.e., $[u,w]^{\top}={\bf R}^{-1}(\theta){\bm{v}}$. The airspeed is $V=\sqrt{u^{2}+w^{2}}$, and the angle of attack is given by $\alpha=\tan^{-1}(-w/u)$  [^6] Here, we abuse the notation $\alpha$ to denote the angle of attack; elsewhere, $\alpha$ denotes an extended class $\mathcal{K}_{\infty}$ function.. We denote the various aerodynamic coefficients as $C_{(\cdot)}\in\mathbb{R}$.

### v4

@@ I INTRODUCTION
[…]
@@ II-A Tangent Cone
Here, $\textup{dist}({\bm{y}},\mathcal{S})$ denotes the distance from a point ${\bm{y}}\in\mathbb{R}^{n}$ to $\mathcal{S}$. If $\mathcal{S}$ is convex, the $\liminf$” can be replaced by a $\lim$” in (3), and $\mathcal{T}_{\mathcal{S}}({\bm{x}})$ remains convex. Moreover, for any ${\bm{x}}\in\textup{Int}(\mathcal{S})$ we have $\mathcal{T}_{\mathcal{S}}({\bm{x}})=\mathbb{R}^{n}$, so that the tangent cone is only non-trivial on the boundary $\partial\mathcal{S}$ of $\mathcal{S}$.
@@ III METHOD OVERVIEW
[…]
@@
Consider the closed-loop system (2) with control inputs ${\bm{u}}\in K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}}(t))$. Suppose that the system is initialized with a locally validated CBF parameter $\tilde{\bm{\alpha}}(t_{0})$ at $t=t_{0}$, and that it is adapted at times $(t_{k})_{k\in\mathbb{N}\cup\{0\}}$ with $t_{k+1}-t_{k}<T$ for all $k$ such that $\tilde{\bm{\alpha}}(t_{k})$ satisfies (22) over $t_{k},t_{k}+T]$. Then, the system trajectories remain within $\mathcal{S}$ for all $t\geq t_{0}$.
@@
We prove this by showing that for all $k\in\mathbb{N}\cup\{0\}$, the state remains in $\mathcal{S}$ over the interval $t_{k},t_{k}+T]$. The proof is by induction.
@@
Inductive Step: Suppose the claim is true for some $k>0$, i.e., the state remains in $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{k}))\subseteq\mathcal{S}$ for all $t\in[t_{k},t_{k}+T]$. By the adaptation, the CBF parameter is updated at some time $t_{k+1}$ such that it is locally validated over $t_{k+1},t_{k+1}+T]$, where $t_{k+1}<t_{k}+T$. Again, by Definition 7 and Nagumo’s Theorem, the system is ensured to remain in $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{k+1}))\subseteq\mathcal{S}$ over $t_{k+1},t_{k+1}+T]$. ∎
@@ V-A Probabilistic Ensemble Neural Network Model
We employ a Probabilistic Ensemble Neural Network (PENN) as the prediction model, denoted as $\bf F$}, to capture both prediction uncertainty and inherent output variability [18]. Building on the input $\bar{X}$ from Sec. IV, we augment it with the CBF parameter of interest $\tilde{\bm{\alpha}}$. Given this augmented input $X=[\bar{X}^{\top},\tilde{\bm{\alpha}}$$^{\top}$, the model outputs two quantities $Y=[\phi,\delta]^{\top}$: a safety margin $\phi\in\mathbb{R}$ and a performance metric $\delta\in\mathbb{R}$ (see Fig. 2b). Concretely, we define $\phi$ as the minimum of the ICCBF constraint function over the interval $[t,t+T]$:
[…]
@@ V-B2 Verification Under Aleatoric Uncertainty
[…]
@@ V-C Online CBF Parameter Adaptation
[…]
@@ VI-A1 Aerodynamics Forces
Let $(u,w)$ be the velocity in the body frame, i.e., $u,w]^{\top}={\bf R}^{-1}(\theta){\bm{v}}$. The airspeed is $V=\sqrt{u^{2}+w^{2}}$, and the angle of attack is given by $\alpha=\tan^{-1}(-w/u)$  [^6] Here, we abuse the notation $\alpha$ to denote the angle of attack; elsewhere, $\alpha$ denotes an extended class $\mathcal{K}_{\infty}$ function.. We denote the various aerodynamic coefficients as $C_{(\cdot)}\in\mathbb{R}$.